OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (31356):

40%

OpenStudy (anonymous):

Hint: less 100% XD

OpenStudy (31356):

90-50= 40%

OpenStudy (31356):

Medal?

OpenStudy (31356):

Lol Sourwing.

OpenStudy (the_fizicx99):

Medals, pfft. Useless. onle virtual =.=

OpenStudy (poopsiedoodle):

Pretty sure it's 80%, since 80% of 50 is 40, and it's the % increase.

OpenStudy (anonymous):

here is better hint. between 50% and 100%

OpenStudy (anonymous):

I agree.^

OpenStudy (poopsiedoodle):

So, since 50+40 is 90, you're adding 80% of 50, which is the % increase.

OpenStudy (31356):

Actually it's 80% sorry messed up on my math.

OpenStudy (31356):

50 to 90 90 - 50 = 40 40/50 = 80% increase

OpenStudy (the_fizicx99):

Medal.. pfffffft

OpenStudy (anonymous):

math can you help me

OpenStudy (anonymous):

yes or no

OpenStudy (anonymous):

no